Required information Problem 13-5A Comparative ratio analysis LO A1, P3 [The following information applies to the questions d1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A Invent Turnover 1A Days Sal in Inv 1A Days Sal Uncol Reg 1B For both1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A Invent Turnover 1A Days Sal in Inv 1A Days Sal Uncol Req 1B For both1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A Invent Turnover 1A Days Sal in Inv 1A Days Sal Uncol Reg 1B For both1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A Invent Turnover 1A Days Sal in Inv 1A Days Sal Uncol Reg 1B For both1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A Invent Turnover 1A Days Sal in Inv 1A Days Sal Uncol Reg 1B For bothComplete this question by entering your answers in the tabs below. 1A Current Ratio 1A Acid Test Ratio 1A Acct Rec Turn 1A In
Answer #1

Answer:

1.Current Ratio = Current Assets / Current Liabilities

Barco = ( 18500 + 38400 + 10100 + 84440 + 5100) / 61340 = 2.55

Kyan = (34000 + 58400 + 8400 + 130500 + 7150 ) / 104300 = 2.29

2.Acid Test Ratio = (Cash + Accounts Receivable + Current Notes Receivable) / Current Liabilities

Barco = ( 18500 + 38400 + 10100 ) / 61340 = 1.09

Kyan = (34000 + 58400 + 8400 ) / 104300 = 0.966

3.Accounts Receivable Turnover = Sales / Avg.Receivables

Barco = 810000 / ( 38400 + 10100 + 28800 + 0 ) / 2 = 810000 / 38650 = 20.96

Kyan = 891200 / ( 58400 + 8400 + 57200 + 0 ) / 2 = 891200 / 62000 = 14.37

4.Inventory Turnover = Cost of goods sold / Avg.Inventory

Barco = 587100 / ( 84440 + 65600 ) / 2 = 587100 / 75020 = 7.83

Kyan = 638500 / ( 130500 + 115400 ) /2 = 638500 / 122950 = 5.19

5.Days Sale in Invenrtory = 365 / Inventory Turnover

Barco = 365 / 7.83 = 46.62 Days

Kyan = 365 / 5.19 = 70.33 Days

6.Days Sales Uncollected = 365 / Accounts Receivable Turnover

Barco = 365 / 20.96 = 17.41 Days

Kyan = 365 / 14.37 = 25.4 Days

Barco has better short term credit risk